OpenStudy (anonymous):

You place $5,000 in a bank account that earns an annual rate of 4.25%. If the interest is compounded quarterly, how much money will you have at the end of 5 years, assuming you neither deposit nor withdraw any money in this time? Select one: a. $6,156.73 b. $6,176.90 c. $6,183.83 d. $6,515.63

OpenStudy (anonymous):

@cwrw238 @eliassaab

OpenStudy (tbnrfrags):

whats \[5,000\div4.25?\]

OpenStudy (anonymous):

1176.47?

OpenStudy (tbnrfrags):

yes, thats 4.25% of 5000 so whats next?

OpenStudy (anonymous):

Well if I multiply it by 5 I don't get any of the multiple choice answers...soooo what do I do next @TBNRfrags

OpenStudy (tbnrfrags):

it's paid out quarterly how often is that?

OpenStudy (anonymous):

4..

OpenStudy (tbnrfrags):

Your money gains interest quarterly in a year....4 times a year. But its for 5 years so 5 * 4 = 20. Your money accrues interest for 20 periods

OpenStudy (anonymous):

So what do I do

OpenStudy (tbnrfrags):

Your money gains interest quarterly in a year....4 times a year. But its for 5 years so 5 * 4 = 20. Your money accrues interest for 20 periods

OpenStudy (anonymous):

Okay so..what now :(

OpenStudy (tbnrfrags):

here i'll tell you the answer then you can try to figure out why thats the answer okay?

OpenStudy (tbnrfrags):

your answer is B

OpenStudy (anonymous):

\[ 5000 \left(1+\frac{0.0425}{4}\right)^{20}=6176 .9 \] B